# Angular integrals with three denominators via IBP, mass reduction, dimensional shift, and differential equations

### v1

@@ Appendix D Proof of the general dimensional-shift identity
In this appendix we prove the general dimensional-shift identity eq. (23). The proof is a straightforward generalization of the two-denominator case given in [5]. We start from the hypergeometric representation of the one-denominator angular integral in $d=4-2\varepsilon$ dimensions,
@@
$I_{j_{1}\dots j_{n}}$ $=\int\frac{\mathrm{d}\Omega_{d-1}}{\Omega_{d-3}}\frac{1}{(v_{1}\cdot k)^{j_{1}}\cdots(v_{n}\cdot k)^{j_{n}}}\,$
@@
$x_{1}^{j_{1}}\cdots x_{n}^{j_{n}}\,\delta\!\left(1-\sum_{i=1}^{n}x_{i}\right)=\left[x_{1}^{j_{1}+1}x_{2}^{j_{2}}\cdots x_{n}^{j_{n}}+x_{1}^{j_{1}}x_{2}^{j_{2}+1}\cdots x_{n}^{j_{n}}+x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{n}^{j_{n}+1}\right]\delta\!\left(1-\sum_{i=1}^{n}x_{i}\right),$ (100)
@@ Appendix E Expansion of dimensionally-shifted two-denominator integrals
In this appendix we list the $\varepsilon$-expansions of the dimensionally-shifted two-denominator integrals used in the differential equation in section VII. They are constructed from the known expansion in $d=4-2\varepsilon$ dimensions [3, 4, 5] via the identities
@@
$I^{(1)}_{1}(v_{11};\varepsilon-1)$ $=\frac{1-2\varepsilon}{2(1-\varepsilon)(1-v_{11}}\Big{[}I^{(0)}(\varepsilon)-v_{11}\,I^{(1)}_{1}(v_{11};\varepsilon)\Big{]}\,,$ (103)
$I^{(0)}_{1,1}(v_{12};\varepsilon-1)$ $=-\frac{1}{2-v_{12}}\left[\frac{1-2\varepsilon}{2\varepsilon}\,I^{(0)}(\varepsilon)+v_{12}\,I^{(0)}_{1,1}(v_{12};\varepsilon)\right],$ (104)
$I^{(1)}_{1,1}(v_{12},v_{11};\varepsilon-1)$ $=-\frac{v_{12}}{v_{12}(2-v_{12})-v_{11}}\left[\frac{1-2\varepsilon}{2\varepsilon}\,I^{(0)}(\varepsilon)-\left(1-\frac{v_{11}}{v_{12}}\right)I^{(1)}_{1}(v_{11};\varepsilon)+v_{12}\,I^{(1)}_{1,1}(v_{12},v_{11};\varepsilon)\right],$ (105)
@@ Appendix F Calculation of the boundary value for the massless master integral
The boundary value for $J^{(0),d=6}_{111}$ is calculated at the symmetric point $v_{12}=v_{13}=v_{23}=1$, where the three vectors are in an orthogonal configuration depicted in figure 2. At this point the Euclidean Gram determinant becomes unity and the defining integral representation eq. (1) becomes
$J^{(0),d=6}_{111}(1,1,1)=I^{(0),d=6}_{111}(1,1,1)=\frac{1}{2}\int_{0}^{\pi}\mathrm{d}\theta_{1}\int_{0}^{\pi}\mathrm{d}\theta_{2}\int_{0}^{\pi}\mathrm{d}\theta_{3}\frac{\sin^{3}\theta_{1}\sin^{2}\theta_{2}\sin\theta_{3}}{(1-\cos\theta)(1-\sin\theta_{1}\cos\theta_{2})(1-\sin\theta_{1}\sin\theta_{2}\cos\theta_{3})}\,.$ (110)
Expanding the third denominator in a geometric series and integrating the $\theta_{3}$-integrals, using that
@@
$J^{(0),d=6}_{111}(1,1,1)=\sum_{n=0}^{\infty}\frac{1}{2n+1}\int_{0}^{\pi}\mathrm{d}\theta_{1}\int_{0}^{\pi}\mathrm{d}\theta_{2}\frac{\sin^{3+2n}\theta_{1}\sin^{2+2n}\theta_{2}}{(1-\cos\theta)(1-\sin\theta_{1}\cos\theta_{2})}=\sum_{n=0}^{\infty}\frac{1}{2n+1}\,I_{1,1}^{(0)}(v_{12}=1;\varepsilon=-n-1)\,,$ (112)
where we could identify the remaining integral with a massless two-denominator angular integral in $6+2n$ dimensions. This integral has been calculated in $d$ dimensions in terms of the Gauss hypergeometric function [2, 4, 5],
@@
$I_{1,1}^{(0)}\!\left(\frac{1}{2};-n-1\right)=\frac{\pi}{n+1}\,{}_{2}\mathrm{F}_{1}\left(1,1,2+n;\frac{1}{2}\right)=\pi\left[\psi^{(0)}\!\left(\frac{n}{2}+1\right)-\psi^{(0)}\!\left(\frac{n}{2}+\frac{1}{2}\right)\right].$ (115)
@@
$J^{(0),d=6}_{111}(1,1,1)=\pi\,\sum_{n=0}^{\infty}\frac{1}{4n+1}\left[\psi^{(0)}\!\left(n+1\right)-\psi^{(0)}\!\left(n+\frac{1}{2}\right)\right]+\pi\,\sum_{n=0}^{\infty}\frac{1}{4n+3}\left[\psi^{(0)}\!\left(n+\frac{3}{2}\right)-\psi^{(0)}\!\left(n+1\right)\right].$ (116)
@@
where $A_{n}=\sum_{k=1}^{n}(-1)^{k+1}/k$ denotes the alternating harmonic series we find
@@
$J^{(0),d=6}_{111}(1,1,1)=2\pi\sum_{n=0}^{\infty}\frac{(-1)^{n}}{2n+1}\left(\log 2-A_{n}\right).$ (123)
@@
$J^{(0),d=6}_{111}(1,1,1)=2\pi\,\int_{0}^{1}\frac{\mathrm{d}t}{t}\,\arctan t=2\pi\,\mathrm{Ti}_{2}(1)=2\pi\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{(2n+1)^{2}}=2\pi\,\mathrm{C}\,,$ (126)
where $\mathrm{Ti}_{2}$ is the inverse tangent integral and $\mathrm{C}$ is Catalan’s constant.
@@ Appendix G Useful integrals for the calculation of the master integrals
From Lewin’s delightful book [53] we have the following integrals useful for the integration of eqs. (53) and (59),
@@
$J_{1}(\alpha,\delta,x)$ $=\int_{-\alpha}^{x}\mathrm{d}t\,\frac{\delta\log|\alpha+t|}{\delta^{2}+t^{2}}=\frac{\theta+\eta}{2}\log\left(\alpha^{2}+\delta^{2}\right)-\frac{1}{2}\mathrm{Cl}_{2}(2\theta+2\eta)-\frac{1}{2}\mathrm{Cl}_{2}(\pi-2\theta)-\frac{1}{2}\mathrm{Cl}_{2}(\pi-2\eta)\,,$ (128)
@@
$+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\theta-2\varphi\right)+\frac{1}{2}\mathrm{Cl}_{2}\!\left(2\theta+2\varphi\right)-\mathrm{Cl}_{2}\!\left(2\theta\right),$ (129)
@@ Appendix H Clausen function
The Clausen function $\mathrm{Cl}_{2}(\theta)$, named after 19th century Danish astronomer and mathematician Thomas Clausen, who calculated the function to 16 decimal places in 1832 [54], is defined by the Fourier series
@@
By differentiating eq. (132) one obtains the integral representation [54, 53]
@@
Figure 4: The left panel shows a plot of the Clausen function on the interval $[0,2\pi]$. The right panel displays the combination $\mathrm{Cl}_{2}(2\arctan x)$ as it appears in the results from section VIII.
@@
which utilizes the quick convergence of the even zeta values $\zeta_{2n}\equiv\sum_{k=1}^{\infty}\frac{1}{k^{2n}}=(-1)^{n-1}\frac{(2\pi)^{2n}B_{2n}}{2(2n)!}$ to $1$ for large $n$. It was this series, truncated at $n=14$, Clausen used in his original calculation [54]. Other efficient numerical implementations can be found in [55, 56, 57].
@@
$\mathrm{Cl}_{2}(2\arctan x)=\mathrm{D}_{2}\!\left(\frac{\mathrm{i}-x}{\mathrm{i}+x}\right)\,\quad\text{respectively}\quad\mathrm{Cl}_{2}(2\arctan x+2\arctan y)=\mathrm{D}_{2}\!\left(\frac{\mathrm{i}-x}{\mathrm{i}+x}\,\frac{\mathrm{i}-y}{\mathrm{i}+y}\right).$ (140)
The Bloch-Wigner dilogarithm is connected to the volume of ideal hyperbolic tetrahedra [58]. Concretely, in the upper-half space model, where the hyperbolic space $\mathfrak{H}_{3}$ is represented by $\mathbb{C}\times\mathbb{R}_{+}$, a tetrahedron with vertices $z_{0,1,2,3}\in\mathbb{C}$ has the hyperbolic volume
@@
$\mathrm{Cl}_{2}(2\arctan(z))=\frac{1}{2\mathrm{i}}\left[\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}-z}{\mathrm{i}+z}\right)-\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}+z}{\mathrm{i}-z}\right)\right]=\mathrm{i}\left(\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}+z}{\mathrm{i}-z}\right)+\frac{1}{4}\log^{2}\!\left(\frac{z+\mathrm{i}}{z-\mathrm{i}}\right)+\frac{\pi^{2}}{12}\right).$ (144)
It has a branch cut along the imaginary axis from $-\mathrm{i}$ to $\mathrm{i}$. In the high energy physics literature the Clausen function and its related constants appear in a variety of Feynman integral calculations, for example in [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 22].
Acknowledgements.

### v4

@@ Appendix D Proof of the general dimensional-shift identity
In this appendix we prove the general dimensional-shift identity eq. (24). The proof is a straightforward generalization of the two-denominator case given in Lyubovitskij:2021 . We start from the hypergeometric representation of the one-denominator angular integral in $d=4-2\varepsilon$ dimensions,
@@
$I_{j_{1},\dots,j_{n}}$ $=\int\frac{\mathrm{d}\Omega_{d-1}}{\Omega_{d-3}}\frac{1}{(v_{1}\cdot k)^{j_{1}}\cdots(v_{n}\cdot k)^{j_{n}}}\,$
@@
$x_{1}^{j_{1}}\cdots x_{n}^{j_{n}}\,\delta\!\left(1-\sum_{i=1}^{n}x_{i}\right)=\left[x_{1}^{j_{1}+1}x_{2}^{j_{2}}\cdots x_{n}^{j_{n}}+\cdots+x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{n}^{j_{n}+1}\right]\delta\!\left(1-\sum_{i=1}^{n}x_{i}\right),$ (101)
@@ Appendix E Expansion of dimensionally-shifted two-denominator integrals
In this appendix we list the $\varepsilon$-expansions of the dimensionally-shifted two-denominator integrals used in the differential equation in section 7. They are constructed from the known expansion in $d=4-2\varepsilon$ dimensions Beenakker:1988 ; Somogyi:2011 ; Lyubovitskij:2021 via the identities
@@
$I^{(1)}_{1}(v_{11};\varepsilon-1)$ $=\frac{1-2\varepsilon}{2(1-\varepsilon)(1-v_{11})}\Big[I^{(0)}(\varepsilon)-v_{11}\,I^{(1)}_{1}(v_{11};\varepsilon)\Big]\,$, (104)
$I^{(0)}_{1,1}(v_{12};\varepsilon-1)$ $=-\frac{1}{2-v_{12}}\left[\frac{1-2\varepsilon}{\varepsilon}\,I^{(0)}(\varepsilon)+v_{12}\,I^{(0)}_{1,1}(v_{12};\varepsilon)\right],$ (105)
$I^{(1)}_{1,1}(v_{12},v_{11};\varepsilon-1)$ $=-\frac{v_{12}}{v_{12}(2-v_{12})-v_{11}}\left[\frac{1-2\varepsilon}{2\varepsilon}\,I^{(0)}(\varepsilon)-\left(1-\frac{v_{11}}{v_{12}}\right)I^{(1)}_{1}(v_{11};\varepsilon)\right$. (106)
$\left.\qquad\qquad\qquad\qquad+v_{12}\,I^{(1)}_{1,1}(v_{12},v_{11};\varepsilon)\right],$
@@ Appendix F Calculation of the boundary value for the massless master integral
The boundary value for $J^{(0),d=6}_{1,1,1}$ is calculated at the symmetric point $v_{12}=v_{13}=v_{23}=1$, where the three vectors are in an orthogonal configuration depicted in figure 2. At this point the Euclidean Gram determinant becomes unity and the defining integral representation eq. (1) becomes
$J^{(0),d=6}_{1,1,1}(1,1,1)=I^{(0),d=6}_{1,1,1}(1,1,1)$
$=\frac{1}{2}\int_{0}^{\pi}\mathrm{d}\theta_{1}\int_{0}^{\pi}\mathrm{d}\theta_{2}\int_{0}^{\pi}\mathrm{d}\theta_{3}\frac{\sin^{3}\theta_{1}\sin^{2}\theta_{2}\sin\theta_{3}}{(1-\cos\theta)(1-\sin\theta_{1}\cos\theta_{2})(1-\sin\theta_{1}\sin\theta_{2}\cos\theta_{3})}\,.$ (111)
Expanding the third denominator in a geometric series and integrating the $\theta_{3}$-integrals [^5] To be mathematically precise, care must be taken whenever interchanging a series with an integration. In this particular case one can convince oneself that the partial sums of the geometric series are – for an odd upper bound – bounded by the non-expanded expression, so it only remains to show that the original integral is finite for Lebesgue’s dominated convergence theorem to be applicable. Finiteness of the original integral can be established by checking that the integrand is finite at the spurious singularities at $\theta_{1}=0$ respectively $\theta_{1}=\pi/2$, $\theta_{2}=0$, and only grows as $r^{-1}$ ($r$ being the distance to the singularity) near $(\theta_{1},\theta_{2},\theta_{3})=(\pi/2,\pi/2,0)$ which is integrable in a $3$ dimensional integral. , using that
@@
$J^{(0),d=6}_{1,1,1}(1,1,1)$ $=\sum_{n=0}^{\infty}\frac{1}{2n+1}\int_{0}^{\pi}\mathrm{d}\theta_{1}\int_{0}^{\pi}\mathrm{d}\theta_{2}\frac{\sin^{3+2n}\theta_{1}\sin^{2+2n}\theta_{2}}{(1-\cos\theta)(1-\sin\theta_{1}\cos\theta_{2})}$
$=\sum_{n=0}^{\infty}\frac{1}{2n+1}\,I_{1,1}^{(0)}(v_{12}=1;\varepsilon=-n-1)\,$, (113)
where we could identify the remaining integral with a massless two-denominator angular integral in $6+2n$ dimensions. This integral has been calculated in $d$ dimensions in terms of the Gauss hypergeometric function vanNeerven:1985 ; Somogyi:2011 ; Lyubovitskij:2021 ,
@@
$I_{1,1}^{(0)}\!\left({\color[rgb]{0,0,0}1};-n-1\right)=\frac{\pi}{n+1}\,{}_{2}\mathrm{F}_{1}\left(1,1,2+n;\frac{1}{2}\right)=\pi\left[\psi^{(0)}\!\left(\frac{n}{2}+1\right)-\psi^{(0)}\!\left(\frac{n}{2}+\frac{1}{2}\right)\right].$ (116)
@@
$J^{(0),d=6}_{1,1,1}(1,1,1)$ $=\pi\,\sum_{n=0}^{\infty}\frac{1}{4n+1}\left[\psi^{(0)}\!\left(n+1\right)-\psi^{(0)}\!\left(n+\frac{1}{2}\right)\right]$
$+\pi\,\sum_{n=0}^{\infty}\frac{1}{4n+3}\left[\psi^{(0)}\!\left(n+\frac{3}{2}\right)-\psi^{(0)}\!\left(n+1\right)\right].$ (117)
@@
where $A_{n}=\sum_{k=1}^{n}(-1)^{k+1}/k$ denotes the alternating harmonic numbers we find
@@
$J^{(0),d=6}_{1,1,1}(1,1,1)=2\pi\sum_{n=0}^{\infty}\frac{(-1)^{n}}{2n+1}\left(\log 2-A_{n}\right).$ (124)
@@
$J^{(0),d=6}_{1,1,1}(1,1,1)=2\pi\,\int_{0}^{1}\frac{\mathrm{d}t}{t}\,\arctan t=2\pi\,\mathrm{Ti}_{2}(1)=2\pi\sum_{n=1}^{\infty}\frac{(-1)^{n+1}}{(2n+1)^{2}}=2\pi\,\mathrm{C}\,$, (127)
where $\mathrm{Ti}_{2}$ is the inverse tangent integral and $\mathrm{C}$ is Catalan’s constant. A direct numerical check of the original integral eq. (111) additionally confirms $J_{1,1,1}^{(0),d=6}(1,1,1)\approx 5.75518$ further raising confidence in the correctness of the result also for physicists who are not inclined to go through the motions akin to footnote 5 in each step.
@@ Appendix G Useful integrals for the calculation of the master integrals
From Lewin’s delightful book Lewin:1981 we have the following integrals useful for the integration of eqs. (54) and (60),
@@
$J_{1}(\alpha,\delta,x)$ $=\int_{-\alpha}^{x}\mathrm{d}t\,\frac{\delta\log|\alpha+t|}{\delta^{2}+t^{2}}=\frac{\theta+\eta}{2}\log\left(\alpha^{2}+\delta^{2}\right)-\frac{1}{2}\mathrm{Cl}_{2}(2\theta+2\eta)$
$\qquad\qquad\qquad\qquad\qquad-\frac{1}{2}\mathrm{Cl}_{2}(\pi-2\theta)-\frac{1}{2}\mathrm{Cl}_{2}(\pi-2\eta)\,$, (129)
@@
[…]
@@ Appendix H Clausen function
The Clausen function $\mathrm{Cl}_{2}(\theta)$, named after 19th century Danish astronomer and mathematician Thomas Clausen, who calculated the function to 16 decimal places in 1832 Clausen:1832 , is defined by the Fourier series
@@
By differentiating eq. (133) one obtains the integral representation Clausen:1832 ; Lewin:1981
@@
Figure 4: The left panel shows a plot of the Clausen function on the interval $0,2\pi]$. The right panel displays the combination $\mathrm{Cl}_{2}(2\arctan x)$ as it appears in the results from section 8.
@@
which utilizes the quick convergence of the even zeta values $\zeta_{2n}\equiv\sum_{k=1}^{\infty}\frac{1}{k^{2n}}=(-1)^{n-1}\frac{(2\pi)^{2n}B_{2n}}{2(2n)!}$ to $1$ for large $n$. It was this series, truncated at $n=14$, Clausen used in his original calculation Clausen:1832 . Other efficient numerical implementations can be found in Wood:1968 ; Kolbig:1995 ; Wu:2010 .
@@
$\mathrm{Cl}_{2}(2\arctan x)=\mathrm{D}_{2}\!\left(\frac{\mathrm{i}-x}{\mathrm{i}+x}\right)\;\;\text{respectively}\;\;\mathrm{Cl}_{2}(2\arctan x+2\arctan y)=\mathrm{D}_{2}\!\left(\frac{\mathrm{i}-x}{\mathrm{i}+x}\,\frac{\mathrm{i}-y}{\mathrm{i}+y}\right).$ (141)
The Bloch-Wigner dilogarithm is connected to the volume of ideal hyperbolic tetrahedra Zagier:2007 . Concretely, in the upper-half space model, where the hyperbolic space $\mathfrak{H}_{3}$ is represented by $\mathbb{C}\times\mathbb{R}_{+}$, a tetrahedron with vertices $z_{0,1,2,3}\in\mathbb{C}$ has the hyperbolic volume
@@
$\mathrm{Cl}_{2}(2\arctan z)$ $=\frac{1}{2\mathrm{i}}\left[\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}-z}{\mathrm{i}+z}\right)-\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}+z}{\mathrm{i}-z}\right)\right]$
$=\mathrm{i}\left(\mathrm{Li}_{2}\!\left(\frac{\mathrm{i}+z}{\mathrm{i}-z}\right)+\frac{1}{4}\log^{2}\!\left(\frac{z+\mathrm{i}}{z-\mathrm{i}}\right)+\frac{\pi^{2}}{12}\right).$ (145)
It has a branch cut along the imaginary axis from $\mathrm{i}$ to $\mathrm{i}$. In the high energy physics literature the Clausen function and its related constants appear in a variety of Feynman integral calculations, for example in Lu:1992 ; Wagner:1996 ; Davydychev:1997 ; Broadhurst:1998 ; Broadhurst:1999 ; Davydychev:2000 ; Davydychev:2003 ; Czakon:2004 ; Binger:2006 ; Adams:2015 ; Chicherin:2020 ; Abreu:2022 ; Devoto:2024 .